### v2

# Biparental Reproduction may Enhance Species Sustainability by Conserving Shared Parental Traits more Faithfully than Monoparental Reproduction

Abstract
Recognized effects of sexual reproduction and other forms of biparental reproduction in species sustainment and evolution include the increasing of diversity, accelerating adaptation, constraining the accumulation of deleterious mutations, and the homogenization of species genotype. Nevertheless, many questions remain open with regard to the evolution of biparental reproduction. In this paper we contribute an initial exploration of finer details of the homogenization effect that we believe deserve focused analysis and discussion. Specifically, rudimentary mathematical analyses suggest the perspective that biparental reproduction enhances the retention in the offspring of shared, i.e., common, properties of the parent generation, as compared with monoparental, clonal reproduction. We argue that this is an intrinsic effect of merging the encodings of parents’ traits, regardless of physical, chemical, biological and social aspects of the reproduction process and of the trait at hand. As the survival of offspring often depends on their ability to step in and actively fill the voids left by failing, dying or absent members of the species in interaction networks that sustain the species, cross-generation retention of common species properties helps sustain the species. This side-effect of sexual reproduction may have also contributed to the very evolution and pervasiveness of sexual reproduction in nature.
1 Introduction
Natural reproduction is of course about transferring parental features to offspring. It is also well accepted that biparental reproduction yields resemblance of offspring in certain aspects to one parent or the other, or to both, or to having a mixture of traits from both parents. In this paper we explore a finer effect of sexual reproduction which is not addressed explicitly in discussions of evolution of sexual reproduction like those cited below, namely, that biparental reproduction conserves traits that are shared by both parents, and that may be common to the entire parent generation within the species.
While there are multiple theories for the sustainment and evolutionary effects of biparental reproduction, both sexual (with two or more mating types) and unisexual (with one mating type, as is common, for example in fungi [Heitman, 2015]), over monoparental, clonal, asexual reproduction, the underlying question is still considered open [MacPherson et al., 2023, MacPherson et al., 2021, Livnat and Papadimitriou, 2016, Billiard et al., 2012]. and others.
The most commonly cited consequences of biparental reproduction include: (i) the introduction of diversity, which then leads to increased adaptivity to changing conditions, and enhanced competitiveness and the ability to co-evolve with pathogens, parasites, or symbionts; for example the Red Queen hypothesis [Van Valen, 1973]; (ii) the constraining the irreversible penetration of deleterious mutations, often termed Muller’s ratchet [Kondrashov, 1988, Muller, 1932]; (iii) preservation of common properties by facilitating the raising of reproduction barriers between divergent subgroups within a species [Barton et al., 2007, CH.22], DNA repair [Bernstein et al., 1981], [Barton et al., 2007, Ch.12 p.335] that is part of meiosis, sexual selection which promotes reproduction of individuals with certain preferred properties, and interference with specialization and speciation of variants that find and prefer new niches [Barton et al., 2007, Ch.16]. Indeed, the genome homogenization effect of biparental reproduction is often taken for granted when a population is panmictic, i.e., every organism of the species can mate with every other of the same species, and confined within a geographical region [Fitzpatrick et al., 2008, Abstract]; it is also often implicit from discussions of gene flow in structured populations [Barton et al., 2007, Ch.16 p.441.], or in pondering the puzzle of sympatric speciation, i.e., why do panmictic populations in a given location sometime give rise to diversity and speciation [Fitzpatrick et al., 2008, Mayr, 1982].
Here we show that in a population whose clonal reproduction is imperfect and may induce some change, as is the case with many kinds of DNA mutations [Barton et al., 2007, Ch.12], horizontal gene transfer [Keeling and Palmer, 2008] incorporation of viral genetic material [Koonin, 2010], etc., biparental reproduction that employs merging of encodings of inherited properties, has an underlying mathematical conservation effect that is independent of biological specifics, and is more general than the homogenization mechanisms listed above. Specifically, biparental reproduction conserves the properties that are common within the parents’ generation better than monoparental reproduction. This, of course, then extends to conservation of ancestral properties across multiple generations.
We support our claim that the effects of biparental reproduction are intrinsic, by initially showing that they exist in an abstracted setting: we deal with only a single property, and the mechanism does not attempt to model in detail meiosis, genetic recombination, epistasis, syngamy, the role of mating types, the role of the environment, and other complex aspects of reproduction in nature. While all these factors are of course important and should be part of realistic models of reproduction in nature, the particular aspects of biparental reproduction discussed here deserve a special focus beyond the published theories, mechanisms and effects.
1.1 Current theories do not imply conservation of shared traits.
Below we list some distinctions between offspring’s retention of traits that are common to their two parents, from presently published theories and observations around sexual reproduction.
  1. 1.
    Consider at first the prominent effect of sexual reproduction, which is the enhancing species diversity. This capability relies on the fact that individuals differ from each other in the first place. By contrast, the effect discussed here focuses on the traits that do not change
  2. 2.
    Plain merging of encodings is not the same as repair: It does not involve a preferred template, and does not call for assessment of whether a structure is broken or intact.
  3. 3.
    Conservation of shared properties is not about elimination of traits that interfere with fitness. In the present model, entity survival is not affected by whether its traits are deleterious or beneficial to its ability to reproduce, or its rate of reproduction.
  4. 4.
    There is no preferential attitude to properties, in contrast, for example, to how sexual selection contributes to species homogenization.
  5. 5.
    The present analysis is different from the homogenization effect of reproductive barrier raised between divergent subspecies, as the group we analyze is the entire population where every individual can and may mate with every other individual in the species.
  6. 6.
    The way sexual reproduction and associated gene flow constrains specialization and fitting into new niches by diluting new beneficial traits may be a special case of the mechanism described here. Biparental reproduction slows down property drift regardless of how useful the new properties may be in new niches. We emphasize that interference with speciation comes from retention of shared traits at the expense of diverse ones. Hypothetically the dilution of over-specialized properties could occur by further randomizing then, and preventing them from becoming common and shared, without slowing down the drifting away of the properties that are shared and common.
  7. 7.
    Most importantly, homogenization of the next generation as induced by sexual reproduction, and discussed extensively in the literature, does not necessarily imply that a shared property of a homogeneous new generation is also closer to the shared property of the parent generation. For example, if the parent generation is comprised of clusters of different trait values, the child generation may have genotype and phenotype properties that are not in any of these clusters.
In the next section, we briefly discuss some of the motivation for the particular principles applied in the ensuing model, and in Section 3 we use a synthetic model with encodings consisting of sets of real numbers to prove mathematically the inevitability of conservation of parental properties as a general effect of biparental reproduction derived from basic probability and the central limit theorem. In future versions we plan to add numerical examples and simulation runs. Version 1 of the arXiv preprint contains such numerical examples, and analysis and visualization of simulation runs of a slightly different model.
In summary, given that survival of organisms, species and entire ecosystems depends on organisms fitting in into a rich interaction network, and given that all living individuals are destined to eventually die, our interim results show intrinsic properties of biparental reproduction likely help sustain each species by enabling new generations to carry out the interaction roles performed by their parents and by several preceding generations. Further studies are needed to generalize and expand the model, and to confirm that the conservation of shared traits by biparental reproduction is indeed a distinct natural phenomenon.
2 Modeling Rationale
Before delving into the details of the mathematical model, in this section we offer some background and context for various choices applied in the model and its analysis. The reader may choose to skip this section, and return to it as needed.
2.1 Codes, Interaction codes and Autoencoding
The present work is part of a larger research project. In [Cohen and Marron, 2023, Cohen and Marron, 2020] and related papers we have outlined how sustainment of a species depends on interaction networks in which it participates. In such networks, the role of each species is naturally and automatically encoded in what we term species interaction code, manifested in the genes, physiology and environment of the individuals of the species. We aim to extend our modeling to address not only species properties, but inter- and intra-species interactions, and research how sexual reproduction may help encode and conserve species interaction codes over multiple generations.
As part of our work on natural autoencoding we have experimented with autoencoding of basic structures like images of straight line segments [Bayer et al., 2023], which hypothetically, can be encoded in just four numbers: the two pairs of coordinates of the endpoints, or a pair of coordinates, an angle and a length. While not dealing with autoencoding per-se, the present model further reduces this two dimensional line encoding problem to a one dimensional continuous range of real numbers which can be encoded in just two numbers, like the value of its boundaries, or its center point and its length. For the formal proof we used a normal distribution with a given mean and variance, rather than a continuous segment.
In this spirit, we focus on the passing down of the encoding of a trait, not the trait itself, even though the phenotype is what in the end sustains the organism. For example, we assume that if the property of interest of an individual is its average height as an adult, or its expected life span, or its varying colors in various seasons, then this property must somehow be encoded before it can be recreated. The encoding itself does not have to be compact or minimalistic; for example, it may involve extensive redundancy, to overcome decoding errors, or may be captured physically, as is the case in a human-made diagram or graph, which may be recorded in many pixels.
This perspective connects the present work to quantitative genetics [Barton et al., 2007, Ch.14;16;17;19;26],[Kang, 2020],[Fisher, 1918]. While blending inheritance theory, where parents traits are mixed to form the offspring features, is long considered obsolete, quantitative genetics studies the many quantitative phenotypic traits, like lifespan, weight, strength, or speed, which are affected by vast genetic information, commonly represented in Quantitative Trait Loci (QTL) which may include multiple genes, regulatory sequences, epigenetic effects, etc. The phenotype aggregation of such polygenic features involves among others operations that resemble adding and averaging the various genetic factors within the organism. Furthermore, the way offspring QTL information is determined from the merging of parents’ QTL is known to be highly complex [Cui et al., 2004, Niu et al., 2022], and cannot yet be reduced to a small set of rules; the model makes a modest attempt to reflect this complexity.
Throughout this paper, we use the term an encoding to refer to an object or structure that encodes, or represents, another object or concept. Thus, an encoding is synonymous here with the term a code, which we presently reserve for the broader concept of species interaction code.
While the model does not aim at mimicking meiosis and creation of gametes, the model includes an elaborate step of recreating a parent’s encoding that will then serve as a seed for subsequent reproduction steps; this is instead of just directly copying the parent’s encoding.
2.2 Information aggregation by set operations
While the relation to quantitative phenotype traits allows some measure of numerical calculations on traits, we wanted the determination of offspring trait encodings to be carried out largely by set operations (like set union) on parent encodings, relying less on pure numerical operations on the two sources.
The present setting is amenable to extensions using more elaborate methods for deriving child traits from parents’
Note that pure averaging of a single property represented by a random variable would be analogous to combining two one-dimensional random walks, with the same step size and equal probabilities for stepping in either direction. Such random walks can be readily shown to have a mean position which is the average of the two origins, and, more importantly, its variance is half the (equal) variance of the two original random walks. This is aligned with the known fact that biparental reproduction reduces the variance within the offspring generation.
2.3 Minimalism, abstraction and accessible language
The model is not intended to be a full model of what happens in nature. In fact, we try to show that in situations of noisy inheritance, it is the very merging of parent encodings that enables the cross-generation conservation of ancestral traits. Examples of elements of the abstraction that we employed include: (i) The property of interest and its encoding both involve real numbers. Extending the arguments to pure symbolic entities, or concrete objects like molecules, is left for a future research; (ii) The inevitable variation in natural reproduction is modeled by replacing direct copying with sampling from a normal distribution; (iii) the biparental reproduction is unisexual, i.e., there are no distinct mating types.
Since here we aim to only show relations between certain values, rather than establishing those values, we adopt the “story proof” argumentation style, as defined and used in[Blitzstein and Hwang, 2015]; this style relies on text of intuitive, yet well founded, logic statements to support the claims, and keeps mathematical formulas to a minimum. In addition, the occasional use of terms that are commonly applied to populations of living organisms is only for convenience.
3 Problem formalization and claims
In this section we compare biparental and monoparental reproduction in the context of a specific model.
3.1 The underlying model
Consider the following synthetic model of a population of entities, also referred to as individuals. The size of the population is kept fixed at a number $n$. The analysis does not include considerations of time. The population reproduces in methods discussed below. A set of $n$ entities, referred to as the cohort of the next, or child, generation, appears instantly and synchronously.
In the present discussion it does not matter whether the parent generation persists or disappears as long as the entities do not reproduce. For simplicity we assume that the entire cohort of each parent generation disappears immediately and synchronously upon the appearance of the cohort of the child generation. One reason for this choice is to avoid dealing with inter-generation mating, in a population in which diversity among co-existing distant generations is unbounded.
We label the individual entities in a cohort as $\{e_{1},e_{2},...,e_{n}\}$;
All entities are associated with exactly one “phenotype” property $P$, which is manifested as a real number representing some quantity. Different entities may have the same or different values for $P$. For illustration, the abstract entities may be thought of as fixed-width sticks where $P$ is their length, or as lumps of material where $P$ is their weight. Let $p$ denote any value of $P$, and, for example, $p_{i}$ is the value of $P$ in the entity $e_{i}$. As for the focus on a quantitative property, see [Barton et al., 2007, Ch.16 p.441]. With each new generation, under the two reproduction methods, the $p_{i}$ values and their distribution within each new cohort may change. In each cohort $c$ we are first interested in the mean of the $p_{i}$ values within the cohort
$p_{c}=\frac{1}{n}\sum_{i=1}^{n}p_{i}~{}~{}.$
In each parent generation $c$, one can view $p_{c}$ as representing a common property, or capability, which is manifested with some diversity in the living individuals. And, if the value of $p$ in an individual is essential for survival within the species’ ecosystem, then the distribution of the property $P$ among the members of the next generation $c^{\prime}$ affects the likelihood that individuals from $c^{\prime}$ will be able to participate in ecological networks in which their parents from $c$ no longer function.
The “genotype” encoding of $P$ in each individual $e_{i}$ is a set $g_{i}$ of an even number $k$ of real values which are also referred to as the points in $g_{i}$. We allow the rare case of any two points in a set being equal to each other, without resorting to other mathematical terms like $g_{i}$ being a multiset or a tuple. The integer $k$ is a fixed parameter in the model, and it is large, say, $k\geq 100$.
The phenotype manifestation of the property $P$ in $e_{i}$ is the mean of the points in $g_{i}$. That is, given an individual $E_{i}$, $1\leq i\leq n$, with genotype encoding $g_{i}=\{x_{i_{1}},x_{i_{2}},...,x_{i_{k}}\}$,its phenotype value $p_{i}$ is computed as follows:
$p_{i}=\frac{1}{k}\sum_{\ell=1}^{k}x_{i_{\ell}}~{}~{}.$
3.2 Reproduction methods
We model and compare the effects of two different reproduction processes, termed Monoparental Reproduction (MR, hereafter) and Biparental Reproduction (BR, hereafter).
Below we describe one reproduction step, termed a transition in each of the methods.
Monoparental Reproduction
  1. 1.
    Denote the current population as being the set $c$ and create an empty set $c^{\prime}$.
  2. 2.
    Repeat $n$ times to create $n$ children:
    1. (a)
      With equal probability, pick a single individual $e_{i}\in c$, by picking a random integer $1\leq i\leq n$; the selected individual will be the single parent of the child created in this iteration; in separate iterations, the same individual $e_{i}$ may be selected again. Note: hereafter, when the distribution of a random choice is not stated, it is assumed to be uniform, with equal probability.
    2. (b)
      Let $g_{i}=\{x_{i_{1}},x_{i_{2}},...,x_{i_{k}}\}$ be the genotype encoding of $e_{i}$. Compute the mean of $g_{i}$, which is equal to the phenotype of $e_{i}$:
      $p_{i}=\frac{1}{k}\sum_{\ell=1}^{k}x_{i_{\ell}}~{}~{};$
      this choice of reproduction step simplifies the model while keeping it intuitive and not limiting the claims; it can be generalized in future analyses.
    3. (c)
      Compute a new set $g^{\prime}$ by sampling $k$ real numbers, $g^{\prime}=\{x^{\prime}_{1},x^{\prime}_{2},...,x^{\prime}_{k}\}$, subject to normal distribution with mean $p_{i}$ and standard deviation $h$ (and variance $h^{2}$) for some fixed global parameter $h$ which specifies and constrains the noise, or error, that may be introduced in the reproduction processes.
    4. (d)
      Create an entity $e^{\prime}$ whose genotype encoding is $g^{\prime}$, and its phenotype is the mean of the points of $g^{\prime}$, and add it to $c^{\prime}$.
  3. 3.
    After completing $n$ iterations: Mark the current population as being the set $c^{\prime}$ and discard $c$.
Biparental Reproduction
  1. 1.
    Denote the current population as being the set $c$ and create an empty set $c^{\prime}$.
  2. 2.
    Repeat $n$ times to create $n$ children:
    1. (a)
      Pick a random pair of two separate individuals, $e_{u},e_{v}\in c,~{}u\neq v,1\leq u,v\leq n$ out of the $\frac{n(n-1)}{2}$ possible pairs in $c$. The pair $e_{u},e_{v}$ will be the joint parents of the child created in this iteration; in separate iterations these two individuals may participate in other pairs or again in the very same pair.
    2. (b)
      Let the respective genotype encodings of $e_{u}$ and $e_{v}$ be
      $g_{u}=\{x_{u_{1}},x_{u_{2}},...,x_{u_{k}}\}$ and
      $g_{v}=\{x_{v_{1}},x_{v_{2}},...,x_{v_{k}}\}$, and let $p_{u}$ and $p_{v}$ be the respective means of these genotype encodings; again, these mean values are conveniently equal to the respective phenotypes.
    3. (c)
      Create two interim (“gamete”) encodings $g^{\prime}_{u}$ and $g^{\prime}_{v}$ by randomly selecting $k/2$ points, under normal distribution with mean $p_{u}$ and variance $h^{2}$, and $k/2$ points under normal distribution with mean $p_{v}$ and variance $h^{2}$, respectively; let $g^{\prime}$ be the union of these two sets, $g^{\prime}=g^{\prime}_{u}\cup g^{\prime}_{v}$. We term this process merging of the parents’ encodings. Note that this merging involves a sampling step, and does not rely on copying of elements of $g_{u}$ or $g_{v}$ into $g^{\prime}$.
    4. (d)
      Create an entity $e^{\prime}$ whose genotype encoding is $g^{\prime}$, and its phenotype is the mean $p^{\prime}$ of the points of $g^{\prime}$, and add it to $c^{\prime}$.
  3. 3.
    After completing $n$ iterations: Mark the current population as being the set $c^{\prime}$ and discard $c$.
As stated earlier, neither MR nor BR employ considerations of Natural Selection and classical fitness metrics like an individual’s probability of reproduction and its expected number of offspring over its life time, and BR does not make use of mating types or of mate selection.
3.3 Analytical claims
In comparing MR and BR in a particular transition from a parent cohort $c$ to a child cohort $c^{\prime}$, we are initially interested in a basic metric $d$ defined as the distance between the phenotype properties of a random individual in $c^{\prime}$, denoted $p^{\prime}_{i}$, and the common property $p_{c}$ of the parents’ cohort
$d_{i}=|p^{\prime}_{i}-p_{c}|$
For a given parent cohort $c$, let $d_{\text{MR}}$ and $d_{\text{BR}}$ be the expected values $E[d_{i}]$ under the MR and BR processes, respectively, when considering all possible $c\shortrightarrow c^{\prime}$ transitions that could emanate from $c$ in one reproduction step, and random choices of $e^{\prime}_{i}$ individuals within $c^{\prime}$.
For a given parent cohort $c$, let $\sigma^{2}_{\text{MR}}$ and $\sigma^{2}_{\text{BR}}$ be the expected values $E[\sigma^{2}_{c^{\prime}}]$, the variance of $c^{\prime}$ under the MR and BR processes, respectively, when considering all possible $c\shortrightarrow c^{\prime}$ transitions that could emanate from $c$ in one reproduction step under the respective reproduction process.
Proposition 1. Let $c$ be a cohort of $n$ entities in the above model; assume that $n$ is large and that the distribution of the $p_{i}$ values of individuals in $c$ approximates a random sample from a normal distribution with mean $p_{c}$ and some variance $\sigma^{2}_{c}$;
Then, $d_{\text{BR}}<d_{\text{MR}}$.
Proof.
A. Compute $d_{\text{MR}}$.
Let $c^{\prime}_{\text{MR}}$ be a random child cohort of $c$, selected from all possible MR transitions with $c$ as a parent.
Let $e^{\prime}_{i_{\text{MR}}}$ be a random individual in $c^{\prime}_{\text{MR}}$.
Let $g^{\prime}_{i_{\text{MR}}}$ be the genotype encoding of $e^{\prime}_{i_{\text{MR}}}$ and let $p^{\prime}_{i_{\text{MR}}}$ be its phenotype.
Let $e_{i}$ be the element of $c$ that served as the single parent of $e^{\prime}_{i_{\text{MR}}}$ in this MR transition.
Let $p_{i}$ be the phenotype of $e_{i}$ (it is also the mean of the genotype encoding $g_{i}$ of $e_{i}$).
The genotype encodings $g^{\prime}_{i_{\text{MR}}}$ of all possible direct MR children of $e_{i}$, can be considered as samples of size $k$ from a normal distribution with mean $p_{i}$ and variance $h^{2}$.
According to the central limit theorem the means of these samples, namely $p^{\prime}_{i_{\text{MR}}}$, are distributed normally with mean $p_{i}$ and standard deviation $\frac{h}{\sqrt{k}}$ (and variance $\frac{h^{2}}{k}$).
We are facing now a compound, nested distribution, that involves (i) the random selection of an $e_{i}$, which is close enough for our purposes to drawing a random $p_{i}$ from the original normal distribution of $c$ with mean $p_{c}$ and variance $\sigma^{2}_{c}$, and (ii) based on the selected $p_{i}$ and the noise parameter, randomly selecting the sample $g^{\prime}$ of size $k$ and computing its mean $p^{\prime}_{i_{\text{MR}}}$.
We can now examine the distribution of the random variable $p^{\prime}_{i_{\text{MR}}}$ as drawn from the original distribution of $c$. According to the law of total variance (a.k.a. Eve’s law) [Blitzstein and Hwang, 2015, Sec.9.5], this is a normal distribution with mean $p_{c}$ and variance that is the sum of the variances, hence
$\sigma^{2}_{c^{\prime}_{\text{MR}}}=\sigma^{2}_{c}+\frac{h^{2}}{k}$
In general, the average absolute deviation of a random sample from the mean of a normal distribution with standard deviation $\sigma$ is $\sigma\cdot\sqrt{\frac{2}{\pi}}~{}~{}$  [Geary, 1935]. Therefore, the expected value of the absolute distance of $p^{\prime}_{i_{\text{MR}}}$ from $p_{c}$ is
$d_{\text{MR}}=E[|p^{\prime}_{i_{\text{MR}}}-p_{c}|]=\left(\sqrt{\sigma^{2}_{c}+\frac{h^{2}}{k}}\right)\cdot\sqrt{\frac{2}{\pi}}~{}~{}.$
B. Compute $d_{\text{BR}}$.
We now examine the expected effects of all possible BR transitions with $c$ as a parent cohort.
Let $c^{\prime}_{\text{BR}}$ be a random child cohort of $c$ within all possible BR transitions.
Let $e^{\prime}_{j_{\text{BR}}}$ be a random individual in $c^{\prime}_{\text{BR}}$.
Let $g^{\prime}_{j_{\text{BR}}}$ be the genotype encoding of $e^{\prime}_{j_{\text{BR}}}$ and and let $p^{\prime}_{j_{\text{BR}}}$ be its phenotype.
Let $e_{u}$ and $e_{v}$ be the two elements of $c$ that served as the parents of $e^{\prime}_{j_{\text{BR}}}$ in this BR transition, and let $p_{u}$ and $p_{v}$ be the means of their genotype encodings, respectively.
The genotype encoding $g^{\prime}_{j_{\text{BR}}}$ was created as a union of two random samples $g^{\prime}_{u}$ and $g^{\prime}_{v}$, each of size $k/2$, around the means $p_{u}$ and $p_{v}$, respectively. Let $p^{\prime}_{u}$ and $p^{\prime}_{v}$ be the means of $g^{\prime}_{u}$ and $g^{\prime}_{v}$, respectively.
The phenotype $p^{\prime}_{j_{\text{BR}}}$ is the mean of the union of $g^{\prime}_{u}$ and $g^{\prime}_{v}$, and is thus the average of the means of these two equal size samples:
$p^{\prime}_{j_{\text{BR}}}=\frac{p^{\prime}_{u}+p^{\prime}_{v}}{2}$
Let us now analyze the random selection of $p^{\prime}_{j_{\text{BR}}}$ compounding the randomness inherent in the way $g^{\prime}_{u}$ and $g^{\prime}_{v}$ were sampled, and the nested randomness of $p_{u}$ and $p_{v}$ as selected from the original distribution of $c$.
The distribution from which both $p_{u}$ and $p_{v}$ were sampled is normal with mean $p_{c}$ and variance $\sigma^{2}_{c}$.
The sets $g^{\prime}_{u}$ and $g^{\prime}_{v}$ were drawn from distributions with means $p_{u}$ and $p_{v}$ respectively, both with variance $h^{2}$. According to the Central Limit Theorem again, the distributions of $p^{\prime}_{u}$ and $p^{\prime}_{v}$ are normal with mean $p_{u}$ and $p_{v}$ respectively, and standard deviation $\frac{h}{\sqrt{\frac{k}{2}}}$ (and variance $\frac{h^{2}}{\frac{k}{2}}$).
Compounding each of these two distributions separately with the underlying distribution of $c$ we get that $p^{\prime}_{u}$ and $p^{\prime}_{v}$ are distributed normally with mean $p_{c}$ and variance $\sigma^{2}_{c}+\frac{h^{2}}{\frac{k}{2}}$
Clearly the mean of the random variable $p^{\prime}_{j_{\text{BR}}}$ is $p_{c}$ since it is half of the sum $E[p^{\prime}_{u}]$ and $E[p^{\prime}_{v}]$, and its variance of is a quarter of the variance of their sum, as it is is the square of the scaling of the variable and the mean:
$\sigma^{2}_{\text{BR}}=\frac{1}{4}\cdot 2(\sigma^{2}_{c}+\frac{h^{2}}{\frac{k}{2}})=\frac{\sigma^{2}_{c}}{2}+\frac{h^{2}}{k}.$
As before,
$d_{\text{BR}}=E[|p^{\prime}_{j_{\text{BR}}}-p_{c}|]=\left(\sqrt{\frac{\sigma^{2}_{c}}{2}+\frac{h^{2}}{k}}\right)\cdot\sqrt{\frac{2}{\pi}}~{}~{}.$
C. Comparing $d_{\text{BR}}$ and $d_{\text{MR}}$.
Thus, $d_{\text{BR}}<d_{\text{MR}}$. $\square$
Note that an intermediate result in the above calculations is that $\sigma^{2}_{\text{BR}}<\sigma^{2}_{\text{MR}}$. This aligns well with the established fact, as was discussed earlier, that in nature, sexual reproduction yields a lower variance within the new generation than does monoparental reproduction. In the present context, such smaller values of $\sigma_{c^{\prime}}$ help individuals in $c^{\prime}$ step in and replace each other when needed, independently of the magnitude of their differences from their parents and from earlier generations.
4 Conclusion and next steps
We have shown that mathematically, biparental, or sexual, reproduction can contribute to conservation in the offspring of traits that are common and important in sustainment of the parent generation. Confirming the observation in nature, and investigating it in more elaborate models, including multiple properties, multiple generations, and more detailed reflection of biological aspects, remains as future work.
Acknowledgements
We thank Netzer Moriya and Hagai Cohen for valuable discussions. This research was funded in part by an NSFC-ISF grant to DH, issued jointly by the National Natural Science Foundation of China (NSFC) and the Israel Science Foundation (ISF grant 3698/21). Additional support was provided by a research grant to DH from Louis J. Lavigne and Nancy Rothman, the Carter Chapman Shreve Family Foundation, Dr. and Mrs. Donald Rivin, and the Estate of Smigel Trust.